OpenStudy (anonymous):

what is the square root of 98?

OpenStudy (anonymous):

9.8994949

jimthompson5910 (jim_thompson5910):

\[\large \sqrt{98}\] \[\large \sqrt{49\times 2}\] \[\large \sqrt{49}\times \sqrt{2}\] \[\large 7\sqrt{2}\] So \[\large \sqrt{98}=7\sqrt{2}\]